OpenStudy (anonymous):

Prove:

OpenStudy (anonymous):

\[(1 + \tan \theta \tan \beta)^{2} + (\tan \theta - \tan \beta)^{2}\] \[= \sec ^{2} \theta \sec ^{2} \beta\]

OpenStudy (anonymous):

Is the question correct, plz recheck it i got some other answer ?

OpenStudy (anonymous):

Yes, its correct. You have to prove that the LHS is equal to the RHS.

OpenStudy (asnaseer):

\[\begin{align} (1 + \tan \theta \tan \beta)^{2} + (\tan \theta - \tan \beta)^{2}&=1+\tan^2(\theta)\tan^2(\beta)+2\tan(\theta)\tan(\beta)\\ &+\tan^2(\theta)-2\tan(\theta)\tan(\beta)+\tan^2(\beta)\\ &=1+\tan^2(\theta)+\tan^2(\beta)+\tan^2(\theta)\tan^2(\beta) \end{align}\]then use the relation:\[\sec^2(\alpha)=\tan^2(\alpha)-1\]in the above and the answer should "pop out" :)

OpenStudy (asnaseer):

sorry, that last relation should be:\[\tan^2(\alpha)=\sec^2(\alpha)-1\]

OpenStudy (anonymous):

Thanks nikhil and asnaseer for the help.

OpenStudy (asnaseer):

yw